How many ions are present in 30.0 mL of 0.600 M Na2CO3 solution?
Answer #1
mol Na2CO3 in 30mL of 0.600M solution
30/1000*0.6 = 0.018 mol Na2CO3
This is 0.018*6.022*10^23 = 1.084*10^22 molecules
Na2CO3 dissociates:
Na2CO3 ? 2Na+ + CO3 2-
1molecule Na2CO3 produces 3 ions
total ions produced = 3*1.084*10^22
Total number of ions = 3.25*10^22 ions in total.
Answer #2

0.600mol/L * 1L/1000mL * 30.0mL = 0.018 moles of Na2CO3.

There are 3 moles of ions (2 Na+ and 1 CO3-) for each moles of Na2CO3 in solution, so triple it!

0.018x3= 0.054moles of ions, which you can multiply by avogadro's number, 6.02x10^23

3.25x10^22 ions